OpenStudy (anonymous):

In establishing standing waves with a Slinky®, your wrist oscillates with f frequency of 8.0 Hz, and the velocity of the wave is 40 m/s. What is the wavelength of this wave? A) 2.5 B) 5.0 C) 10. D) 320

OpenStudy (anonymous):

D. 320. Multiply 8.0 x 40 and you get 320

OpenStudy (aaronq):

for any wave: \(\sf \large wavelength*frequency=velocity \) \(\sf \large wavelength=\dfrac{velocity}{frequency}=\dfrac{40~m/s}{8~/s}=5~m\)